# Recursive Gaussian Process State Space Model

### v3

@@ 4 Adaptation of Inducing points and Hyperparameters
While the recursive learning algorithm for GPSSMs has been developed, two issues remain: the computational burden problem due to the non-parameteric nature of GP, and the hyperparameter online adaptation problem. In Section 4.1, a inducing point set adjustment algorithm is introduced to maintain limited computational complexity, and in Section 4.2, an online hyperparameter adaptation method is derived to improve learning accuracy.
4.1 Dynamic Adjusting the Inducing Points Set
As discussed in Section 3, the computational cost of the learning algorithm increases with the addition of inducing points. This challenge is also encountered in the development of kernel-based online regression algorithms, such as Kernel Recursive Least Squares (KRLS) [51] and Sparse Online Gaussian Processes (SOGP) [27, 52]. These two methods are equivalent to some extent and control computational cost by limiting the size of the inducing point set to a predefined budget $M$. Specifically, this strategy is achieved through two operations: discarding the least important point when the set exceeds the budget $M$, and adding only sufficiently novel points. To implement the discarding and adding operations, the following two key problems must be addressed: (1) how to remove a point with minimal accuracy loss, and (2) how to evaluate the importance or novelty of a point. In the following, we will address these two problems in the context of GPSSMs and develop an adjustment method for the inducing point set, thereby overcoming the computational burden issue. For convenience, let $u_{\mathrm{d}}$ denote the inducing point to be discarded, with index $i_{\mathrm{d}}$ in the point set, and let $\bm{u}_{\mathrm{l}}$ represent the inducing points left.
For the first problem, we define that, after discarding point $u_{\mathrm{d}}$, the new joint distribution $\hat{q}(\bm{x},\bm{f})$, still retains the factorized form (7), namely,
@@ 4.1 Dynamic Adjusting the Inducing Points Set
$\hat{q}^{*}(\bm{x},\bm{u}_{\mathrm{l}})$ $=\arg\min\limits_{\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})}\mathrm{KL}\left[q(\bm{x},\bm{f})\|\hat{q}(\bm{x},\bm{f})\right]$ (20)
@@
Furthermore, for the second problem, the importance or novelty of a point can be assessed by the accuracy loss incurred from optimally discarding it. Specifically, this can be done by calculating the KL divergence in (20) with $\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})=\hat{q}^{*}(\bm{x},\bm{u}_{\mathrm{l}})$:
@@
Through the KL divergence $D^{*}$, the accuracy loss from discarding the inducing point $u_{\mathrm{d}}$ can be evaluated. A computable score $s_{d}$ quantifying the accuracy loss can be obtained by dropping the $u_{\mathrm{d}}$-irrelevant terms in (21) (the derivation is deferred to Appendix 7.1), namely:
@@
where we denote the inverse of the joint covariance as $\bm{\Omega}=\bm{\Sigma}_{t}^{-1}$, and $\bm{\Omega}_{dd}$ is the element of the matrix $\bm{\Omega}$ corresponding to the discarded point $u_{\mathrm{d}}$, namely, its $(d_{x}+i_{\mathrm{d}})$-th diagonal element. In addition, $\bm{Q}_{dd}$ and $\bm{Q}_{du}$ are the $i_{\mathrm{d}}$-th diagonal element and the $i_{\mathrm{d}}$-th row of the inverse kernel matrix $\bm{Q}=\bm{K}_{uu}^{-1}$. Among the terms in (22), $\Delta_{1}$ represents the accuracy loss in the mean of the joint distribution $q(\bm{x}_{t},\bm{f})$, which matches the result in KRLS [51], while $\Delta_{2}$ and $\Delta_{3}$ correspond to the loss in the covariance, which achieve more precise quantification and were not derived in KRLS [51]. Overall, the score $s_{d}$ quantifies the relative accuracy loss from discarding $u_{\mathrm{d}}$, with a smaller score indicating a smaller accuracy loss.
Through (20) and (22), we have addressed the two key challenges. Based on these, we can provide the adjustment rule for the inducing point set, which is divided into discarding and adding operations. For the discarding operation, the rule is as follows: if the size of the inducing point set exceeds the budget $M$, evaluate the score of each point using (22) and remove the one with the lowest score. According to (20), the moments of the new joint distribution $\hat{q}^{*}(\bm{x},\bm{u}_{\mathrm{l}})$ can be easily evaluated by deleting the corresponding elements associated with $u_{\mathrm{d}}$ from the original moments.
For the adding operation, as in the KRLS [51], we will use a more intuitive metric for selecting points, which is derived from the score in (22). Specifically, by observing the score $s_{d}$ in (22), it can be seen that if ${\bm{Q}_{dd}}^{-1}\to 0$, the score will approach its minimum value, i.e., $s_{d}\to$-$\infty$, indicating that the accuracy loss from deleting $u_{\mathrm{d}}$ approaches 0. According to the matrix inversion formula (see Appendix A of [52]), ${\bm{Q}_{dd}}^{-1}=\bm{K}_{dd}-\bm{K}_{dl}\bm{K}_{ll}^{-1}\bm{K}_{ld}$, where subscripts $d$ and $l$ corresponding to $u_{\mathrm{d}}$ and $\bm{u}_{\mathrm{l}}$, and $\bm{Q}_{dd}^{-1}$ actually is the GP prior conditional variance of $u_{\mathrm{d}}$ given $\bm{u}_{\mathrm{l}}$ according to (8). Therefore, we can use the GP prior conditional variance of $f_{t}$ given the current inducing points $\bm{u}$ as the metric for novelty, namely:
@@
where subscript $t$ denotes $f_{t}$. Then, if $\gamma$ is less than a certain threshold, denoted as $\varepsilon_{\text{tol}}$, we will not add the new point to the inducing point set. This filtering process helps to slow down the increase in the size of the inducing point set. Besides, more importantly, it plays a crucial role in ensuring the numerical stability of the learning algorithm. Specifically, when $\gamma$ is small, adding a new point can cause the updated kernel matrix $\bm{K}_{\bar{u}\bar{u}}$ to approach singularity, which would negatively affect the stability of the algorithm. Therefore, the value of the threshold $\varepsilon_{\text{tol}}$ for adding a point can be determined based on the machine accuracy. Accordingly, when not adding points, the moment updating equation (17) will be modificated. We derive this equation in Appendex 7.2, whose result is;
@@
In summary, this subsection derives the optimal deletion method for inducing points and presents a metric for quantifying the importance or novelty of points. Based on these, a dynamic adjustment algorithm, including discarding and adding operations for the inducing point set, is developed to ensure limited computational cost and algorithm stability. In the next subsection, we will address the online hyperparameter optimization problem to enhance learning accuracy.
@@ 4.2 Online Optimization of GP Hyperparameters
In GP learning, if the hyperparameters $\bm{\theta}$ significantly mismatch the function to be learned, it will result in substantial learning errors. In offline learning, the hyperparameters $\bm{\theta}$ can be optimized using data by maximizing the log marginal likelihood [15]. However, as illustrated in the Introduction, the online setting presents two challenges for hyperparameter optimization: 1) the information source problem due to the lack of data retention, and 2) the coupling between GP hyperparameters and the posterior distribution. In this subsection, we will address these two challenges and propose an online GP hyperparameter optimization method for GPSSMs. For clarity, we denote the quantities corresponding to before and after hyperparameter optimization with the superscripts or subscripts "old" and "new", respectively. Additionally, to highlight the impact of the GP hyperparameters, probability distributions depending on $\bm{\theta}$ will be rewritten as $p(\cdot;\bm{\theta})$.
Firstly, to find the information source for hyperparameters optimization, we can first observe the expression of the excat joint posterior distribution, which is given as follows for time $t$:
@@
where $p(\bm{f};\bm{\theta})$ denotes the GP prior on function values $\bm{f}$, and the likelihood model $p(\bm{x}_{t},\bm{y}_{1:t}|\bm{f})=\int\prod_{i=1}^{t}p(\bm{y}_{i},\bm{x}_{i}|\bm{x}_{i-1},\bm{f})p(\bm{x}_{0})\,\mathrm{d}\bm{x}_{0:t-1}$. By using this likelihood model, we can evaluate the marginal likelihood $p(\bm{y}_{1:t};\bm{\theta})=\int p(\bm{f};\bm{\theta})p(\bm{x}_{t},\bm{y}_{1:t}|\bm{f})\,\mathrm{d}\bm{x}_{t}\,\mathrm{d}\bm{f}$, and then optimize the GP hyperparameters by maximizing it. However, for online hyperparameter optimization, an explicit likelihood model is unavailable because the past measurements $\bm{y}_{1:t-1}$ are not stored. Instead, we only have an approximation of the left-hand side of (25), namely the approximate distribution corresponding to the old hyperparameters $\bm{\theta}_{\mathrm{old}}$:
@@
In other words, the measurement information is distilled into the approximate distribution $q_{old}(\bm{x}_{t},\bm{f})$ and the latter can recover the likelihood model, thereby addressing the information source problem.
Secondly, utilizing the information source in (27), we can find the optimization objective for GP hyperparameters and evaluate the posterior distribution after hyperparameter optimization. The key idea in achieving this is to evaluate the joint probability model $p(\bm{f};\bm{\theta})p(\bm{x}_{t},\bm{y}_{1:t}|\bm{f})$, which is proportional to the posterior distribution, and its integral is the marginal likelihood that can serve as the objective function. For convenience, we actually derive the posterior distribution first, which can be obtained by combining the expression of the posterior distribution (25) with the approximate likelihood model (27), namely:
@@
where $q_{\text{new}}(\bm{x}_{t},\bm{f})$ denotes the approximate joint distribution corresponding to the new hyperparameters $\bm{\theta}_{\text{new}}$. In (28), since the distributions $p(\bm{f};\bm{\theta}_{\text{old}})$, $p(\bm{f};\bm{\theta}_{\text{new}})$, and $q_{\text{old}}(\bm{x}_{t},\bm{f})$ are all Gaussian, the new approximate distribution $q_{\text{new}}(\bm{x}_{t},\bm{f})$ is also Gaussian. By expanding the distribution related to $\bm{f}$ in the third row of (28) with the conditional distribution $p(\bm{f}|\bm{u})$, it can be shown that the new distribution retains the factorized form $q_{\text{new}}(\bm{x}_{t},\bm{f})=p(\bm{f}_{\neq u}|\bm{u};\bm{\theta}_{\text{new}})q_{\text{new}}(\bm{x}_{t},\bm{u})$, where:
@@
Therefore, we can obtain the new posterior distribution after the GP hyperparameters change.
Furthermore, by integrating both sides of the first row of (29), we can obtain the marginal likelihood for hyperparameter optimization, which leads to:
@@
Based on these results, the practical optimization objective and moment matching equation for the posterior distribution can be derived, which are presented below.
First, by eliminating some $\bm{\theta}_{\text{new}}$-irrelevant terms in (30) (see Appendix 7.3), a practical optimization objective for GP hyperparameters is to minimize:
@@
Second, the moments of the new posterior $q_{\text{new}}(\bm{x}_{t},\bm{u})$ after the hyperparameter update can be evaluated using a Kalman filter-like equation (derivation in Appendix 7.3):
$\tilde{\bm{H}}=\begin{bmatrix}\bm{0}&\bm{I}_{n_{u}}\end{bmatrix}$ (32)
@@
Remark 4.1
The online optimization method for GP hyperparameters most closely related to this paper is [29], an online GPR method. In this method, the likelihood model is similarly recovered from the approximate posterior distribution, and it is used to optimize the location of inducing points and GP hyperparameters. Clearly, in GPSSMs, it is possible to optimize the inducing inputs in the same way. However, we do not pursue this approach because simultaneously optimizing the inducing inputs and hyperparameters would complicate the loss function and the adjustment equation for the posterior distribution. Additionally, we observe that in GPSSMs, due to the system states being implicitly measured, the difficulty of optimizing the inducing inputs increases. Therefore, we separate the optimization of inducing inputs and GP hyperparameters and determine the inducing-point set using a discretized selection method.

### v4

@@ 4 Adaptation of Inducing points and Hyperparameters
While the recursive learning algorithm for GPSSMs has been developed, two issues remain: the computational burden caused by continuously adding inducing points, and the online hyperparameter adaptation problem. In Section 4.1, an inducing-point set adjustment algorithm is introduced to maintain limited computational complexity, and in Section 4.2, an online hyperparameter adaptation method is derived to improve learning accuracy.
4.1 Dynamically Adjusting the Inducing-Point Set
As discussed in Section 3, the computational cost of the learning algorithm increases with the addition of inducing points. This challenge is also encountered in the development of kernel-based online regression algorithms, such as Kernel Recursive Least Squares (KRLS) [37] and Sparse Online Gaussian Processes (SOGP) [38, 39]. These two methods are equivalent to some extent and control computational cost by limiting the size of the inducing-point set to a predefined budget $M$. Specifically, this strategy is achieved through two operations: discarding the least important point when the set exceeds the budget $M$, and adding only sufficiently novel points. To implement the discarding and adding operations, the following two key problems must be addressed: (1) how to remove a point with minimal accuracy loss, and (2) how to evaluate the importance or novelty of a point. In the following, we will address these two problems in the context of GPSSMs and develop an adjustment method for the inducing-point set, thereby overcoming the computational burden issue. For convenience, let $u_{\mathrm{d}}$ denote the inducing point to be discarded, with index $i_{\mathrm{d}}$ in the point set, and let $\bm{u}_{\mathrm{l}}$ represent the inducing points left.
Optimal deletion for an inducing point. For the first problem, we define that, after discarding point $u_{\mathrm{d}}$, the new joint distribution $\hat{q}(\bm{x},\bm{f})$, still retains the factorized form (7), namely,
@@ 4.1 Dynamically Adjusting the Inducing-Point Set
$\hat{q}^{*}(\bm{x},\bm{u}_{\mathrm{l}})$ (20)
$=\arg\min\limits_{\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})}\mathrm{KL}\left[q(\bm{x},\bm{f})\|\hat{q}(\bm{x},\bm{f})\right]$
@@
Criterion for selecting the least important inducing point. Furthermore, for the second problem, the importance or novelty of a point can be assessed by the accuracy loss incurred from optimally discarding this point. Specifically, this is done by calculating the KL divergence in the third row of (20) with $\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})=\hat{q}^{*}(\bm{x},\bm{u}_{\mathrm{l}})$. The result is:
@@
where $D^{*}$ quantifies the accuracy loss that results from discarding the inducing point $u_{\mathrm{d}}$.
Based on $D^{*}$, we can obtain a computable score to quantify the importance of inducing points. Specifically, we expand the integration in (21) and retain only the terms involving $u_{\mathrm{d}}$, which yields the following score (for the derivation, see Appendix A):
@@
where we denote the inverse of the joint covariance as $\bm{\Omega}=\bm{\Sigma}_{t}^{-1}$, and $\bm{\Omega}_{dd}$ is the element of the matrix $\bm{\Omega}$ corresponding to the discarded point $u_{\mathrm{d}}$, namely, its $(d_{x}+i_{\mathrm{d}})$-th diagonal element. In addition, $\bm{Q}_{dd}$ and $\bm{Q}_{du}$ are the $i_{\mathrm{d}}$-th diagonal element and the $i_{\mathrm{d}}$-th row of the inverse kernel matrix $\bm{Q}=\bm{K}_{uu}^{-1}$. Among the terms in (22), $\Delta_{1}$ represents the accuracy loss in the mean of the joint distribution $q(\bm{x}_{t},\bm{f})$, while $\Delta_{2}$ and $\Delta_{3}$ correspond to the loss in the covariance. As shown in (22), $\Delta_{3}$ involves the joint covariance of $(\bm{x}_{t},\bm{u})$, i.e., $\bm{\Sigma}_{t}$. Therefore, this score can take into account the correlation between the state $\bm{x}$ and the GP, and thus properly quantifies the accuracy loss in the context of GPSSMs. Overall, the score $s_{d}$ quantifies the relative accuracy loss from discarding $u_{\mathrm{d}}$. A smaller score $s_{d}$ indicates a smaller accuracy loss. Therefore, it can be used as the criterion for selecting the least important point to discard.
Through (20) and (22), we have addressed the two main challenges for the selection of the inducing points. Based on these, we can provide the adjustment rule for the inducing-point set, which is divided into discarding and adding operations.
Rule for discarding inducing points. For the discarding operation, the rule is as follows: if the size of the inducing-point set exceeds the budget $M$, evaluate the score of each point using (22) and remove the one with the lowest score. According to (20), the moments of the new joint distribution $\hat{q}^{*}(\bm{x},\bm{u}_{\mathrm{l}})$ can be easily evaluated by deleting the corresponding elements associated with $u_{\mathrm{d}}$ from the original moments.
Rule for adding inducing points. For the adding operation, we use a more convenient and intuitive metric for selecting points, which is derived from the score in (22). Specifically, by observing the score $s_{d}$ in (22), it can be seen that if $\bm{Q}_{dd}^{-1}\to 0$, the score will approach its minimum value, i.e., $s_{d}\to$-$\infty$, indicating that the accuracy loss from deleting $u_{\mathrm{d}}$ approaches 0. According to the matrix inversion formula (see Appendix A of [39]), $\bm{Q}_{dd}^{-1}=\bm{K}_{dd}-\bm{K}_{dl}\bm{K}_{ll}^{-1}\bm{K}_{ld}$, where subscripts $d$ and $l$ correspond to $u_{\mathrm{d}}$ and $\bm{u}_{\mathrm{l}}$. In other words, $\bm{Q}_{dd}^{-1}$ is the GP prior conditional variance of $u_{\mathrm{d}}$ given $\bm{u}_{\mathrm{l}}$ according to (2). Therefore, we can use the GP prior conditional variance of $f_{t}$ given the current inducing points $\bm{u}$ as the metric for novelty, namely:
@@
where subscript $t$ denotes $f_{t}$. Then, if $\gamma$ is less than a certain threshold, denoted as $\varepsilon_{\text{tol}}$, we will not add the new point to the inducing-point set. Through this adding rule, we obtain two advantages. First, this filtering process helps to slow the increase in the size of the inducing-point set. Second, it plays a crucial role in ensuring the numerical stability of the learning algorithm. Specifically, when $\gamma$ is small, adding a new point can cause the updated kernel matrix $\bm{K}_{\bar{u}\bar{u}}$ to approach singularity, which would negatively affect the stability of the algorithm. Therefore, the value of the threshold $\varepsilon_{\text{tol}}$ for adding a point can be determined based on the machine accuracy.
In the above adding rule, there exist cases where points are not added. In this situation, the moment updating equation (17) is no longer applicable. Therefore, we re-derive the moment updating equation for this scenario. The derivation details are provided in Appendix B and the result is as follows:
@@
In summary, this subsection derives the optimal deletion method for inducing points and presents a metric for quantifying the importance or novelty of points. Based on these, a dynamic adjustment algorithm, including discarding and adding operations for the inducing-point set, is developed to ensure limited computational cost and algorithm stability. In the next subsection, we will address the online hyperparameter optimization problem to enhance learning accuracy.
@@ 4.2 Online Optimization of GP Hyperparameters
In GP learning, if the hyperparameters $\bm{\theta}$ significantly mismatch the function to be learned, it will result in substantial learning errors. In offline learning, the hyperparameters $\bm{\theta}$ can be optimized using data by maximizing the log marginal likelihood [14]. However, as illustrated in the Introduction, the online setting presents two challenges for hyperparameter optimization: 1) the information source problem due to the lack of data retention, and 2) the coupling between GP hyperparameters and the posterior distribution. In this subsection, we will address these two challenges and propose an online GP hyperparameter optimization method for GPSSMs. For clarity, we denote the quantities corresponding to before and after hyperparameter optimization with the superscripts or subscripts ”old” and ”new”, respectively. Additionally, to highlight the impact of the GP hyperparameters, probability distributions depending on $\bm{\theta}$ will be rewritten as $p(\cdot;\bm{\theta})$.
Information source for online hyperparameters optimization. Firstly, to find the information source for hyperparameter optimization, we can first observe the expression of the exact joint posterior distribution, which is given as follows for time $t$:
@@
where $p(\bm{f};\bm{\theta})$ denotes the GP prior on function values $\bm{f}$, and the likelihood model $p(\bm{x}_{t},\bm{y}_{1:t}|\bm{f})=\int\prod_{\tau=1}^{t}p(\bm{y}_{\tau},\bm{x}_{\tau}|\bm{x}_{\tau-1},\bm{f})p(\bm{x}_{0})\,\mathrm{d}\bm{x}_{0:t-1}$. By using this likelihood model, we can evaluate the marginal likelihood $p(\bm{y}_{1:t};\bm{\theta})=\int p(\bm{f};\bm{\theta})p(\bm{x}_{t},\bm{y}_{1:t}|\bm{f})\,\mathrm{d}\bm{x}_{t}\,\mathrm{d}\bm{f}$. Then, the GP hyperparameters $\bm{\theta}$ can be optimized by maximizing the marginal likelihood $p(\bm{y}_{1:t},\bm{\theta})$. However, for online hyperparameter optimization, an explicit likelihood model is unavailable because the past measurements $\bm{y}_{1:t-1}$ are not stored. Instead, we only have an approximation of the left-hand side of (25), namely the approximate distribution corresponding to the old hyperparameters $\bm{\theta}_{\mathrm{old}}$:
@@
In other words, the measurement information is distilled into the approximate posterior $q_{\mathrm{old}}(\bm{x}_{t},\bm{f})$ and the latter can recover the likelihood model, thereby addressing the information source problem.
Posterior update equation. Secondly, by utilizing the likelihood model obtained in (27), we can find the optimization objective for GP hyperparameters and evaluate the posterior distribution after hyperparameter update. The key idea in achieving this is to evaluate the joint probability model $p(\bm{f};\bm{\theta})p(\bm{x}_{t},\bm{y}_{1:t}|\bm{f})$, which is proportional to the posterior distribution, and its integration is the marginal likelihood that can serve as the objective function. For convenience, we derive the posterior update equation first, which can be obtained by combining the expression of the posterior distribution (25) with the approximate likelihood model (27), namely:
@@
where $q_{\text{new}}(\bm{x}_{t},\bm{f})$ denotes the approximate joint distribution corresponding to the new hyperparameters $\bm{\theta}_{\text{new}}$. In (28), since the distributions $p(\bm{f};\bm{\theta}_{\text{old}})$, $p(\bm{f};\bm{\theta}_{\text{new}})$ and $q_{\text{old}}(\bm{x}_{t},\bm{f})$ are all Gaussian, the new approximate distribution $q_{\text{new}}(\bm{x}_{t},\bm{f})$ is also Gaussian. By expanding the distribution related to $\bm{f}$ in the third row of (28) with the conditional distribution $p(\bm{f}_{\neq u}|\bm{u})$, it can be shown that the new distribution retains the factorized form $q_{\text{new}}(\bm{x}_{t},\bm{f})=p(\bm{f}_{\neq u}|\bm{u};\bm{\theta}_{\text{new}})q_{\text{new}}(\bm{x}_{t},\bm{u})$, where:
@@
Therefore, we can obtain the new posterior distribution after the GP hyperparameters change. In practical implementation, this can be achieved by a Kalman filter-like equation (the derivation is provided in Appendix C):
$\tilde{\bm{H}}=\begin{bmatrix}\bm{0}&\bm{I}_{n_{u}}\end{bmatrix}$ (30)
Optimization objective. Furthermore, by integrating both sides of the first row of (29), we can obtain the marginal likelihood for hyperparameter optimization, which leads to:
@@
Then, by expanding the probability density in (31) and eliminating terms that are irrelevant to $\bm{\theta}_{\text{new}}$ (the derivation is provided in Appendix C), a practical optimization objective for GP hyperparameters is to minimize:
@@
Remark 4.1.
The online optimization method for GP hyperparameters most closely related to this paper is [40], an online GPR method. In this method, the likelihood model is similarly recovered from the approximate posterior distribution, and it is used to optimize the location of inducing points and GP hyperparameters. Clearly, in the context of GPSSMs, it is possible to optimize the inducing inputs in the same way. However, we do not pursue this approach because simultaneously optimizing the inducing inputs and hyperparameters would complicate the loss function and the posterior update equation. Additionally, we observe that in GPSSMs, due to the system states being implicitly measured, the difficulty of optimizing the inducing inputs increases. Therefore, we separate the optimization of inducing inputs and GP hyperparameters and determine the inducing-point set using a discretized selection method.